A sample of size n = 10 n = 10 is drawn from a population. The data is shown below. 70.4 56.8 39 63.8 70.4 43.8 59.4 46.5 50.4 70.4 What is the range of this data set? range = What is the standard deviation of this data set? (Remember, it is a sample.) Please report the answer with appropriate rounding, reporting 2 more decimal places than the original data. Please, please, please do not calculate the value by hand. stdev =
Answer #1

SOLUTION

Part (a)

Range = maximum value – minimum value

              = 70.4 - 39

              = 31.4 Answer 1

Part (b)

Standard deviation = 12.463 ANSWER 2

Excel Calculations are given below:

i

xi

xi - avg

(xi-avg)^2

1

70.4

13.31

177.1561

2

56.8

-0.29

0.0841

3


39

-18.09

327.2481

4

63.8

6.71

45.0241

5

70.4

13.31

177.1561

6

43.8

-13.29

176.6241

7

59.4

2.31

5.3361

8

46.5

-10.59

112.14181

9

50.4

-6.69

44.7561

10

70.4

13.31

177.1561

Total

570.9

0

1242.6827

Average

57.09

Sample variance =

1242.6827/8

155.335

Sample standard deviation =

sqrt(155.335)

12.463